A decrease in price from R50 to R40 causes the quantity demanded to increase from 250p to 3000 units. Using the arc elasticity of demand, the price elasticity of demand is...
1
Expert's answer
2018-02-22T09:07:08-0500
If a decrease in price from R50 to R40 causes the quantity demanded to increase from 2500 to 3000 units, then using the arc elasticity of demand, the price elasticity of demand is:
Ed = (3000 - 2500)/(40 - 50)*(40 + 50)/(3000 + 2500) = -500/10*90/5500 = -9/11, so the demand is inelastic.